OpenStudy (anonymous):

Lisa spent $67 for a skateboard and $25 for a protective helmet. What percent of the total amount that Lisa spent was the price of the skateboard? Round to the nearest tenth of a percent, if necessary. A. 62.7% B. 72.8% C. 75% D. 81.7%

OpenStudy (anonymous):

@billj5

OpenStudy (anonymous):

@anwaarullah

OpenStudy (anonymous):

Um that's what we have to find out :/

OpenStudy (anwaarullah):

add both

OpenStudy (anonymous):

92

OpenStudy (anwaarullah):

so price of skatboar divide by 92 into 100

OpenStudy (anonymous):

0.92

OpenStudy (anonymous):

To find the percentage you will have to divide the price of the skateboard 67 by the total 92. 67/92 = 72.8%

OpenStudy (anwaarullah):

no divide 67/92*100

OpenStudy (anwaarullah):

when you go direct for answer you wont able to learn ever,,,i want you to understan otherwise i can solve in one sttep

OpenStudy (anonymous):

72.8

OpenStudy (anwaarullah):

nearsrest ten means it would be 73%

OpenStudy (anonymous):

Oh right

OpenStudy (anonymous):

But the choices says 72.8

OpenStudy (anonymous):

Nearest tenth is the first number after the decimal

OpenStudy (texaschic101):

92 is total amount skateboard is 67 what % of total amount was skateboard x% * 92 = 67 x% = 67/92 x% = 0.7282 x = 72.82 rounded = 72.8

OpenStudy (anonymous):

So its b

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

Ok how do I give medals btw

OpenStudy (texaschic101):

whoever you think deserves a medal, click best response

OpenStudy (anonymous):

Oh did I just give you one?

OpenStudy (texaschic101):

It shows I got one...but it is not telling me who gave it to me

OpenStudy (anonymous):

I did lol

OpenStudy (texaschic101):

thanks :)

OpenStudy (anonymous):

Np :)